Two masses of 5kg and 2kg are connected by a rope.


the mass experiences a kinetic friction of 9.8N between it and the table. detemine the accelarition of the system. tension in the rope

1
Expert's answer
2021-11-07T19:29:35-0500

If the 5 kg mass is on the table and the 3 kg mass is hanging connected to the 5 kg mass by a rope over a pulley, we can write the following equations:


"5\text{ kg:}\space Ma=-f+T\\
2\text{ kg:}\space -ma=T-mg,\\\space\\
a=\frac{mg-f}{m+M}=\frac{2·9.8-9.8}{2+5}
=1.4\text{ m/s}^2,\\\space\\
T=m(g-a)=16.8\text{ N}."